r has been placed is 6 weeks What is the reorder point considering only the average demand for the item? 諸units 48 units O 54 units O 66 unitsAnswer to both questions.
Answer #1

Weekly demand d = 8 units

Lead time L = 6 weeks

Reorder point = demand during lead time = dL = 8*6 = 48 units

Reorder point = 48 units

2

Standard deviation \small \sigma = 2 units

Z = 1.65

Safety stock SS

\small SS= \sigma *Z*\sqrt{L}

\small SS= 2*1.65*\sqrt{6}

SS = 8 units